OpenStudy (anonymous):

PLEASE HELP: FAN/MEDAL?!?! If f(x) = log3 (x + 1), what is f−1(2)?

OpenStudy (anonymous):

\[f(x) = \log _{3} (x + 1)\]

OpenStudy (anonymous):

what is \[f ^{-1}(2)\]

OpenStudy (anonymous):

I don't know how to do the inverse of the log....

hartnn (hartnn):

here's a shortcut, find x from \(\Large 2= \log_3(x+1)\)

OpenStudy (anonymous):

ok... so \[2 \div(\log _{3}) -1\] maybe?

hartnn (hartnn):

here's how that works, \(y= \log_3(x+1) \\ x+1 = 3^y \\ x= 3^y-1\) and so, \(f^{-1}(x) = 3^x-1\) do you know the definition of log ? \(\Large \log_ab=c \implies b=a^c\)

OpenStudy (anonymous):

ohh.... ok... so would it be 1?

hartnn (hartnn):

\(\Large 2= \log_3(x+1) \\ \large \implies 3^2 = x+1\) you may want to try to find again from this?

hartnn (hartnn):

try to find "x" again ***

OpenStudy (anonymous):

so 3^2 = x + 1 3^2 = 9 9 = x + 1 8 = x?

hartnn (hartnn):

YES! thats your f^-1 (2) also :)

OpenStudy (anonymous):

Thank you so so so much! :D

hartnn (hartnn):

most welcome ^_^